### v1

# Scalable Radar-based ITS: Self-localization and Occupancy Heat Map for Traffic Analysis

Abstract
4D mmWave radar sensors are well suited for city scale Intelligent Transportation Systems (ITS) given their long sensing range, weatherproof functionality, simple mechanical design, and low manufacturing cost. In this paper, we investigate radar-based ITS for scalable traffic analysis. Localization of these radar sensors in a city scale range is a fundamental task in ITS. For mobile ITS setups it requires more endeavor. To address this task, we propose a self-localization approach that matches two descriptions of "road": the one from the geometry of the motion trajectories of cumulatively observed vehicles, and the other one from the aerial laser scan. An ICP (iterative closest point) algorithm is used to register the motion trajectory into the road section of the laser scan to estimate the sensor pose. We evaluates the results and show that it outperforms other map-based radar localization methods, especially for the orientation estimation. Beyond the localization result, we project radar sensor data onto city scale laser scan and generate an scalable occupancy heat map as a traffic analysis tool. This is demonstrated using two radar sensors monitoring an urban area in the real world.
I Introduction
The transportation system of the future will be supported by Intelligent Transportation Systems (ITS). ITS consist of roadside sensors with different data modalities and the ability to communicate their understanding of the scene to road users. Its constant presence and non-obstructive mounting position can assist vehicles with perception tasks, contributing to safer and more efficient individual travel. It can also provide statistical traffic analysis to support urban plannings. Among the sensors in ITS systems, 4D automotive radar has great potential for smart city applications due to its relatively simple mechanical design and low manufacturing cost. It can provide a 3D point cloud with a radial velocity attribute over a long range under adverse weather conditions. To enhence the work of radar-based ITS, more data from the real world should be collected and used as training data for the development of algorithms. Localization, in terms of determining the installed position and orientation of the sensors, plays a fundamental role for further data processing. This is usually a one-time job with great efforts during the installation phase of test beds with the help of GNSS (Global Navigation Satelite System) technology . For mobile sensor setups (e.g. [1, 2]), which aims at flexibly enriching scenarios, it becomes a more frequent endeavor.
In this work, we investigate scalable radar-based ITS (Fig. 1) by addressing the problem of the localization of ITS radar sensors following by a occupancy heat map generated in a sensor network. More specifically, our localization doesn’t require extra calibration efforts in the sensor installation phase, but directly uses sensor data thereafter. City scale aerial laser scan point clouds are used to aid this process, which are usually available from the local authorities. The pose of the sensor in a global reference frame can be found by registering the 4D radar sensor data into the aerial laser scan. It other words, two descriptions of the concept of "road" are matched together: the one from aerial laser scans, the other one from the motion trajectories of cumulatively observed vehicles. This approach runs online. The result can be checked and corrected on a regular base. Based on this localization result, a further contribution of this work is traffic analysis in the form of occupancy heat maps using multiple sensors for a scalable area. This involves the occupancy of finely subdivided lanelets, featuring both fine-grained and tractable analysis. We demonstrate the flexible deployment of radar-based ITS systems and online generated occupancy heat maps in a scalable manner in the real world.
The rest of this paper is organized as follows: In the second section, we review the related work in sensor localization and traffic analysis visualization. The third section is dedicated to the localization method. Then, in section 4, we provide the evaluation of our localization method. The fifth section presents the occupancy heat map using two sensors in a real-world application. The last section concludes the paper and presents ideas for future research.
Figure 1: Overview of the scalable radar-based ITS. (a) Flexibly deployable sensor setups. (b) Self-localization based on sensor data and aerial laser scan data. The top part of the figure illustrates the processing steps for representing the road section based on the aerial laser scan using a vectorized map. The lower part of the figure illustrates the description of the road section using the radar point cloud data. The voxelized road section from the aerial laser scan and the accumulated radar data are the input to the ICP registration algorithm. (c) Projection of radar sensor data into city scale laser scan. (d) Occupancy heat map as a tool for traffic analysis. Darker parts indicate high utility rate of lanelet sector.
II Related Works
II-A Localization of sensors
Given the increasing popularity of radar sensors, many localization works have been proposed, e.g. [3, 4]. The localization problem is closely related to the mapping process and can therefore be divided into map-based approaches and simultaneous localization and mapping (SLAM). Since ITS are usually static and have a very close relationship to Geographical Information Systems (GIS), map-based approaches fit the purpose well. Open Street Map (OSM) is a community-supported, free world map project. Integrating OSM into the localization task has been a popular research direction. Pioneering works such as [5] and [6] propose to use OSM for camera and lidar localization. Hong et al. [7] combine road routing graph and semantic features in OSM to improve the localization accuracy of radar-only approaches. On a macro scale, OSM provides accurate data for routing and associated localization. However, if the micro-geometry accuracy of the local map within the view of the sensors is not guaranteed, which is usually the case, OSM is not suitable for ITS localization.
The authors of the [8, 9] propose the use of aerial imagery for the localization task of range-finding sensors, both radar and lidar. The overhead images showing the perfect geometry of the road are widely available today from satellites or from survey flights. They use these images to help learn radar images along with their rotation and translation. The aerial laser scans are also increasingly provided by local authorities. This allows exploration of their use for the localization problem. While this approach hasn’t been seen yet, the use of lidar data for radar localization is being intensively explored. [10] learns fake lidar point clouds from radar data for localization. A non-learning approach [11] is proposed to localize the radar-based odometry to a lidar map by finding and matching corresponding locations in both modalities. Most of the work above works with localization of a 2D radar sensor on a moving base. A localization approach for 3D radar in ITS systems is yet to be seen.
Since we use aerial laser scan along with maps for the localization task, the power of HD maps is harvested for high accuracy. Beyond the standard OpenDrive format, we further exploit lane level information. Lanelet2 [12, 13], developed by Poggenhans et al., is an open-source library and framework for detailed high-precision maps based on the concept of Lanelet [14]. A lanelet is the smallest unit within which traffic rules do not change. Furthermore, the topology formed by the lanelets is constant. As a comprehensive and flexible map format, Lanelet/Lanelet2 has recently been widely adopted in ITS tasks (e.g. [15]-[17]). The CommonRoad project [18], which aims to provide various standardized motion planning scenarios, also uses a lanelet-based map format. It further provides a powerful toolkit [19] that allows easy conversion between many different map formats.
II-B Traffic Analysis and ITS Data Visualizations
The long-range sensing capability of radar enables powerful traffic analysis with wide coverage at low equipment count and cost. Even though, city scale traffic analysis needs methods and tools to efficiently fuse the data and generated insights for downsteam tasks [20]. The line chart, the bar chart, the heat map, and the geospatial map are the popular techniques for visualization [21]. The authors in [22] use heat maps for traffic analysis, focusing on spatiotemporal aggregation and distillation. In [23], regional fingerprinting based on heat maps is introduced for monitoring traffic. Another example is VAIT [24], where the heat map is used for visual analysis for urban transportation. We adopt the use of heat map for occupancy in current situation as a base for further traffic analysis.
III Localization
The use of a GNSS-based localization approach for ITS systems has severe limitations. Due to the clearance profile of road design, urban ITS sensors usually have to be installed near trees and buildings, so the multipath and non-line-of-sight (NLOS) effects could negatively influence the accuracy of the GNSS system. In addition, the deployment of a flexible sensor system should also consider the influence of cloudy or rainy weather, where the GNSS receiver is unable to obtain reasonable readings. The data from the radar sensor, on the other hand, shows strong geometric features in the long run given the fixed road structure. After a long enough sampling period, the observed vehicles could pass through any accessible point in the scene. Furthermore, the data is available regardless the weather conditions.
III-A Method Overview
The approach is briefly illustrated in Fig. 1(b). The result of the data flow through two preprocessing pipelines is fed to the ICP algorithm [25]. In the upper part of the Fig. 1(b), the aerial image is vectorized to obtain the HD map. This step is currently performed manually, but could potentially be extended to an automatic process similar to [26]. It is important to preserve the geometric features of the road, so we use the Lanelet2 [12] format for lane-level detail. The map is then projected onto the laser scan frame to mask the portion of the road from the entire scan. The point cloud of the road then goes through the outliner elimination process (e.g. DBSCAN [27]) to get rid of noise such as reflections from trees. The final step of the preprocessing is to voxelize the point cloud into 0.5m $\times$ 0.5m grids. Two versions of the point cloud are used. One is the cloud of grid center points. The other is the data flattened to the minimum height value as a 2D point cloud. This completes the generation of the target point clouds in the ICP algorithm. For real use cases, data from local authorities (e.g. [28]) can be used. In simulations, road meshes can be converted to point clouds for this step.
The second part of the preprocessing is the generation of motion trajectories from radar sensors representing the road section. The first step is to accumulate radar sweeps. A sufficient number of frames $n_{f}$ should be chosen. We set $n_{f}$ to 2000. Since the radar works with 20Hz, we could check and correct the pose of the radar every 100s. Given that the radar could provide about 500 points per scan, a total of 1 million points could be collected, which is too much for the registration algorithm to run efficiently. Since our goal is to retrieve the points that represent the road and the points there should be moving in the long run, static reflections are filtered out. This is done using the Doppler measurement, only points with a radial speed above $0.15m/s$ are considered. A statistical analysis shows that more than $93\%$ of the points have a speed below this threshold, making this criterion an efficient filter that could later speed up the ICP process. The point cloud then goes through an outliner removal process. Ghost reflections and small clusters of points are removed using DBCAN. With the proper parameters, the largest cluster represents the points on the road. The cluster is then voxelized and flattened as source point clouds for the ICP algorithm. This pipeline is shown in the lower part of the Fig. 1(b).
III-B Localization with ICP
The ICP algorithm establishes the data association between the source and target point clouds via nearest neighbor, and reduces the distance between them to obtain the transformation matrix [25]. In our implementation, it follows a multi-scale approach and is executed four times. To initialize the ICP algorithm, we make a manual coarse selection of the deployment position, such as a point on the map, and the estimated orientation, such as "east", "north", etc. The initial ICP process is executed with a large allowed correspondence distance. The following 2 ICP runs use a reduced correspondence distance of 2$\times$ - 1$\times$ voxel size. The source and target input for these runs are 3D point clouds, whereas the last one utilizes the flattened representation of road sections for optimal height matching.
Since the ICP algorithm may encounter a local minimum, we assess the centroid of the corresponding road sections from both the flattened radar point cloud and the flattened aerial laser scan. If the threshold
$\epsilon=\sum_{n}\bm{p}_{n,radar}/n-\sum_{m}\bm{p}_{m,aerial\_laser\_scan}/m$
is exceeded, we translate the radar points to mitigate the error.
For continuous calculation, a Kalman filter with a static motion model is implemented. The state vector $\bm{x}$ consists of the position $\bm{p}$ as a 3D coordinate in the UTM frame with a predefined origin and a quaternion $\bm{q}$ for the orientation.
$\bm{x}=[\bm{p},\bm{q}]^{T}=[x_{p},y_{p},z_{p},x_{q},y_{q},z_{q},w_{q}]^{T}$
The setup presumes a static model. The prediction can be simplified as follows
$\bm{\hat{x}}_{n+1,n}=\bm{\hat{x}}_{n,n}+\bm{w}_{n}$
$\bm{P}_{n+1,n}=\bm{P}_{n,n}+\bm{Q}_{n}$
where $\bm{Q}_{n}$ is the uncertainty of the static model. We utilize the outcome from the ICP algorithm to correct the prediction to obtain the estimate. As ICP generates all values in the state vector, we can simplify the measurement model to $\bm{z}_{n}=\bm{x}_{n}+\bm{v}_{n}$. The correction procedure is given by
$\bm{K}_{n}=\bm{P}_{n,n-1}(\bm{P}_{n,n-1}+\bm{R}_{n})^{-1}$
$\hat{\bm{x}}_{n,n}=\hat{\bm{x}}_{n,n-1}+\bm{K}_{n}(\bm{z}_{n}-\hat{\bm{x}}_{n,n-1})$
$\bm{P}_{n,n}=(\bm{I}-\bm{K}_{n})\bm{P}_{n,n-1}(\bm{I}-\bm{K}_{n})^{T}+\bm{K}_{n}\bm{R}_{n}\bm{K}_{n}^{T}$
where $\bm{R}_{n}$ is the uncertainty of the measurement. The Kalman filter in conjunction with ICP, operates on a 5-second cycle. During each cycle, the radar point clouds are collected to create vehicle traces and perform localization. We implement the localization process on a rolling basis, with the rolling window of 2000 frames. This method filters out temporary changes in sensor pose caused by environmental factors such as wind and rain drops, while preserving long-term changes.
Figure 2: The qualitative evaluation of the localization result is shown in the left figure. Dark blue points are the radar point clouds. Blue to green points are aerial laser scan points. The aligned dark blue clusters exactly overlap the tree canopies (marked with the rectangels). The right image from Google Earth is used as a reference for readers to understand the scene.
IV Evaluation
We evaluate the localization algorithm with both real world data and simulation data.
IV-A Localization of Radar Sensors in Real World
For the real world localization, radar data is gathered in Ingolstadt, Germany with a mobile sensor system [1]. It contains a Continental ARS548 long range radar sensor and a Nvidia Orin embedded system to gather and process the data. The radar spans the azimuth range of 120° and the elevation range of 30° and operates at 20Hz. Each frame delivers about 500 detections in a point cloud format. The aerial laser scan is retrieved from open data provided by the Agency for Digitisation, High-Speed Internet and Surveying of Bavaria [28].
A qualitative assessment is presented in Fig. 2. The localization accuracy is demonstrated using the tree canopies as a reference on a parking lot. The radar data was gathered from 23,950 frames over a period of approximately 20 minutes on a windy day. The point clouds first pass through a velocity filter. Only points with a speed higher than $0.15m/s$ are kept. Subsequently, the DBSCAN algorithm is used to filter the cloud with ($\epsilon=0.5,n_{min}=10$). The points at the canopies of trees represent the reflection from the rustling leaves. It is evident that the points are grouped in clusters with a clear separation between each other. The clusters overlap very well with the laser scan point clouds of the trees. This shows that the algorithm accurately determines the position and orientation.
Figure 3: Quantitative evaluation of the localization. The hexagon is the GNSS measured sensor position. 10 points are randomly selected around the ground truth value of the sensor setup as the initial position for the ICP. The error between the ICP result and the ground truth is given in the table.
For a quantitative evaluation, we place the sensor setup at a random position around a junction to gather data in order to perform the localization. Then we compare the results with a ground truth, which is determined with a GNSS device (ArduSimple 2B) using RTK (real-time kinematics) correction. The accuracy of the GNSS measurement is $2cm$. 10 positions in the map surrounding the test spot as initial values for the ICP algorithm. The actual sensor location and the initial starting points are marked in the Fig. 3. We show the 2D errors within the image. The test runs have similar height outcomes, with an average error of 1.81m. Although the lowest 2D position error is $36cm$, the average position error is 1.08m, which is comparable to other map-based approaches [7, 8, 11]. However, we have observed that the point may move far away from its initial position, even when it is located near the actual ground truth. This is likely due to a hidden pattern in the point cloud data. The altitude error may be due to changes in terrain elevation. In Fig. 4 we see that the vehicle’s tracks align with the driving lane while also leaving the green area, road divider, and other areas unoccupied (upper part). It further depicts that the left-turn vehicles stay correctly in the lane (bottom part). This indicates a acceptable accuracy for traffic analysis.
Figure 4: Alignment of the radar point cloud to the road. The lane description is created manually based on an aerial image (dashed lines). The lane information is also used for the masking process in Fig. 1(b). The upper zoomed-in image shows that the green area on the road matches the non-traveled area in the trajectory point cloud. The lower image shows that the traces of the left-turning vehicles match the lane description.
IV-B Localization of Radar Sensors in Carla Simulation
We evaluate the localization method in Carla simulator’s Town 10 map. A laser scan for this map is not directly available. Instead, we convert the road mesh into a point cloud. We then crop the road section using the HD map of the town to remove the parking lanes. We conduct two sets of experiments that change the location and direction of the radar sensors. In the first set of tests, we place the radar sensor with different yaw angles (DY). In the second set of tests, sensors are located at different positions and have a more flexible orientation (DO). In both sets of tests, the sensors are placed as shown in Fig. 5 The sensors have a maximum sensing range of 200m. Like the radar in the real world, the azimuth angle of the sensors covers from -60° to 60°. Their orientations in the first set are also shown in Fig. 5. The elevation range is 30°. The frame rate is 44Hz.
Figure 5: Test points in the Carla map Town 10. The points and the line segments show the position and orientation of the radar sensors in test DY. In test DO the position is maintained but the orientation is more flexible and not shown in the figure.
Table I shows the result of the experiments. The 2D location error is less than 1 meter, except for R2 and R6 in DO. The altitude estimate is always off by about 1.06 meters on average. The average yaw errors are 0.38° (DY) and 0.75° (DO). In the most favorable scenario, it could only be off by 0.02° (R1 in DY). The roll and pitch errors in DO are 0.28° and 0.33°, respectively. This outperforms other map-based methods [7, 8, 11]. We believe that the favorable outcome of Carla’s 2D position and orientation is due to its complex observable geometry and simple road geometry. The evaluation demonstrates the potential of our approach.
TABLE I: Result of tests in carla
DY Error X(m) Y(m) Z(m) yaw(°)
R1 -0.71 0.68 -1.02 0.02
R2 -0.79 -0.67 -1.05 0.06
R3 0.59 -0.69 -1.05 0.11
R4 -0.09 -0.10 -1.06 0.30
R5 -0.27 0.03 -1.03 0.25
R6 -0.95 -0.69 -1.13 1.53
DO Error XY(m) roll(°) pitch(°) yaw(°)
R1 0.50 0.21 0.12 0.58
R2 5.06 0.16 0.29 0.06
R3 0.59 0.15 0.24 0.16
R4 0.19 0.64 0.41 0.34
R5 0.20 0.19 0.49 0.95
R6 3.35 0.31 0.41 2.42
V Scalable Dynamic Occupancy Map
Figure 6: The global occupancy heat map as a tool for traffic analysis. The heat map is created with data from two sensors. Sensor 1 monitors the road section from top to bottom. Sensor 2 monitors the left-right road section. The maximum range is about 300 meters. The intersection can be monitored by both sensors. The time difference is indicated below the figures. The dark color indicates a high occupancy. The number above the color scale is the maximum occupancy of the polygon nodes in the time scope. For example, 225 implies the occupancy of 11.25s out of 100s ($225\times 50$ms).
After locating multiple radar-based ITS setups in the city (Fig. 1(c)), large-scale traffic analysis could be conducted. It performs better than sensors installed at specific points, such as inductive circuits or ultrasound detectors, within the range. In comparison to vision-based systems, the radar-based systems generate less data traffic while still being able to monitor larger areas using fewer sensors. Using an occupancy heat map as a tool, we can represent the current traffic status and perform spatial statistical analyses.
V-A Using the HD Map as Scaffold
We design the occupancy heat map based on a sublane level HD map. We define the urban occupancy heat map as $M=\{P\}$, where $P$ are the nodes of locations. Unlike scene representation with voxelization, each node in this map is defined as a sub-lane level tile of the road. It consists of a four-vertex polygon with a width of approximately 3m, equal to the street’s width and a length of roughly 0.5m, close to a pedestrian’s standard footstep. These shapes come from a HD map of lanes in Lanelet2 (OSM) format. First, we interpolate the left and right edges of the lanelet to get finer points. Then, every four points that enclose a smallest rectangle form a finer lanelet and are saved as a polygon. A unique ID is assigned to these polygon nodes as a reference. Similar to lanelets at junctions, it is possible to have overlapping nodes.
V-B Local Sensor Data Processing
To analyze the occupancy of the scene, we first map the sensor data to the polygons. Whenever the radar sensor delivers a point cloud, we check their footprints on the east-north plane and assign them to the corresponding polygons. Then, we add the current time stamp to the polygon node’s "occupancy time" attribute, indicating that the node is occupied at that time stamp. Because there may be a very large number of potentially occupied polygons, the R-tree spatial indexing [29] is used for efficient assignment. This compresses the local sensor data into a list of indices of occupied polygons, which can greatly reduce the data volume. Note that the data can be filtered based on the Doppler measurement, the road clearance profile or the radar cross section measurement (reflection strength).
V-C Global Occupancy Heat Map
After connecting sensors to a network, we can generate a global occupancy heat map by fusing all data frames into a single scene representation. We first use NTP (Network-Time-Protocol) to synchronize the time of all data processing units. A 5G module is used to connect the sensor and its data processing unit with the internet. The sensors operate at a nearly constant frequency. However, fusing data from sensors into an integrated map is not as trivial as adding up all the lists of occupied polygons and marking the occupancy with a time stamp. The sensors’ observation areas sometimes overlap and add redundancy. As a result, the overlap areas receive more frequent observations and consequently exhibit higher occupancy. This occurrence suggests incorrectly more occupants in the corresponding areas. To prevent issues, it is necessary to establish a uniform update rate for the global occupancy map that matches the radar update rate. Depending on whether the global occupancy map is generated in a field data processing unit or in the cloud, the update rate can either adopt the frame rate of the connected sensor or be defined independently. Despite efforts to synchronize clocks in data processing units, there is a noticeable time offset in sensor data depending on their starting time (see Fig. 7).
Figure 7: A diagram of message arrival time and time synchronization in the occupancy map. X axis shows the time stamp when messages arrive. Y axis shows the time step between two adjacent messages: $t_{n}=t_{n-1}+y$. The blue rectangles represents sampling windows of 50ms.
The data updates locally at a frequency of approximately 20Hz, but deviations may occur based on the time that onboard algorithms in the sensor may consume. The data for generating the global occupancy heat map is then collected within each 50ms time window. Fig. 7(a) illustrates that if the frame rate is close to 20Hz, we can collect the data and build an occupancy map within each sampling window. If the data frequency deviates significantly (see Fig. 7(b)), two messages may arrive during the same sampling period, resulting in both messages being added to that time step. If no message is received during a particular time step, it should be left empty.
The heat map of occupancy can either be instantly generated to show the current road occupancy status or be cumulatively generated by counting the reported occupancy time stamps to indicate the statistical congestion status. Data volume can be controlled by the number of polygons. Fig. 6 shows an example where two sensors are deployed in the city and localized using the proposed method. At the junction, traffic changes are detected from left to right depending on the occupancy status. Congestion can be observed on the right side of the map, which is due to an increased volume of traffic split into the right turn lane. Similarly, the left-turn waiting area at the intersection from below is congested, indicating a high volume of traffic moving from the upper to the lower portion of the area.
VI Conclusion
In this work, we investigate 4D radar-based ITS to leverage the benefits of long sensing range under diverse weather condition. A method for self-localization is proposed. It leverages idea of matching the concept of "road" in two data modalities: aerial laser scan and live radar sensor data. On one hand, road sections are represented by a point cloud from an aerial laser scan, which is extracted with the help of aerial images and HD maps. On the other hand, the radar point cloud of motion trajectories generated by the observed vehicles also describes the geometry of road sections. These two point clouds are matched with ICP algorithm. The output transformation matrix delivers solutions to the localization task. A extra Kalman Filter can be integrated into the localization process, making corrections during deployments possible. Our results outperforms the result from other map-based radar localization works, especially for the orientation. Using the localization results, we demonstrate the creation of global occupancy heat maps by synchronizing networked radar-based sensor setups for scalable traffic analysis. Based on this work, we contribute to the flexible deployment of radar-based ITS for traffic analysis. The further work lies in detecting individual moving objects on the road and matching their trajectories with the corresponding lane to increase the localization accuracy. Another expansion is adding object information to the occupancy heat map for further traffic analysis.